OpenStudy (anonymous):

Find the x-intercepts of the function. f(x)=x^1/2+1x^1/4-90

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

set f(x) to zero then solve

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

proceed... how do you do that?

OpenStudy (anonymous):

you will need to factor to solve

OpenStudy (anonymous):

keep going...

OpenStudy (anonymous):

oh wait fraction exponents my bad

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

dooh... i thought i was talking to the asker....

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

we can do the same as if we were solving a quadratic..

OpenStudy (anonymous):

i don't understand what to do

OpenStudy (anonymous):

if we let u=x^1/4. so now your equation is in quadratic form: u^2 + u - 90 = 0

OpenStudy (anonymous):

the first thing is x^(1/2) tho ?

OpenStudy (anonymous):

can you solve that equation i just wrote?

OpenStudy (anonymous):

oh wait lol wow sorry

OpenStudy (anonymous):

(1/4)*2 haha =)

OpenStudy (anonymous):

or you can just solve 0 = x^{1/2}+x

OpenStudy (anonymous):

i don't understand how to solve that.

OpenStudy (anonymous):

first of all what is the equation?

OpenStudy (anonymous):

this equation?: u^2 + u - 90 = 0 factor...

OpenStudy (anonymous):

no i'm wondering if this equation is \[x^{1/2}\] or if its a fraction or something lol

OpenStudy (anonymous):

i dont know, i really dont know any of this. i have a list of answers that i can choose from that i can list

OpenStudy (anonymous):

dpanic basically gave you the answer all you have to do is remember FOIL.

OpenStudy (anonymous):

whats is foil?

OpenStudy (anonymous):

okay Ill give an example if you had x^2 + 2x +1 you could also write that as (x+1)(x+1) you just have to follow this same process but for your equation or if that doesn't work use quadratic formula

OpenStudy (anonymous):

so x^2 = (x+1) (x+1)?

OpenStudy (anonymous):

no expand what you see and brackets and you will see it does not equal x^2

OpenStudy (anonymous):

do you know how to expand what you have in brackets?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x^1/2+1x^1/4-90 = 0 Let X = x^1/4, ==> X^2 + X - 90 = 0 Do you understand up to this point?

OpenStudy (anonymous):

Chlorophyll she/he needs to learn factoring in general just ordinary factoring dpanic already got him/her to that point btw

OpenStudy (anonymous):

i dont really understand what u did

OpenStudy (anonymous):

@pleasecrystal that's the whole point why you're on OS website! Take a close look and learn how to solve, will you :)